Circular No. 7997

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ha IN NGC 6962 J. Graham, M. Papenkova, and W. D. Li, University of California

at Berkeley, report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on Oct. 21.2 (mag ∼ 17.3) and 22.2 UT (mag ∼ 16.7). SN 2002ha is located at α = 20^{h}47^{m}18^{s}.58, δ = +0^{o}18′45′′.6 (equinox 2000.0), which is 6′′.9 west and 29′′.4 south of the nucleus of NGC 6962. A KAIT image taken on Oct. 19.2 showed nothing at this position (limiting mag ∼ 19.0).

SAX J1808.4–3658 M. P. Rupen, V. Dhawan, and A. J. Mioduszewski, National Ra-

dio Astronomy Observatory; B. W. Stappers, University of Amsterdam and Stichting Astronomisch Onderzoek in Nederland; and B. M. Gaensler, Harvard-Smithsonian Center for Astrophysics, report the detection of radio emission associated with the recent outburst of the accreting millisecond pulsar SAX J1808.4–3658 (IAUC 7993): “Observations at 8.5 GHz with the Very Large Array (VLA) on Oct. 16.05 UT show a 0.44-mJy source (7σ) at α = 18^{h}08^{m}27^{s}.60 ± 0^{s}.02, δ = −36^{o}58′43′′.4 ± 0′′.4 (equinox 2000.0), consistent with the position reported by Gaensler et al. (1999, Ap.J. 522, L117) during the source’s 1998 outburst. Data taken with the Australia Telescope Compact Array (ATCA) on Oct. 16.07–16.36 give a 3σ detection of 0.4 mJy at 8.5 GHz, suggesting that the source was not rapidly fading. VLA and ATCA observations at 4.9 GHz at the same epochs as above both give a marginal detection at a flux density of 0.3 ± 0.1 mJy, while ATCA observations on Oct. 16.05–16.38 provide an upper limit of 0.5 mJy (3σ) at 2.4 GHz at the detected position. Subsequent VLA observations on Oct. 17.98 at 8.5 GHz give a probable detection (based on positional agreement with the Oct. 16 data) at 0.3 mJy (4σ). Further radio observations are planned; we encourage continued optical and x-ray monitoring.”

CORRIGENDA On IAUC 7996, ‘Possible Nova in Small Magellanic Cloud’, line 2, for

Oct. 19.131 read Oct. 18.131

On IAUC On IAUC 7763, ‘Supernovae’, line 7, for SN 2000fr, the α should read 14^{h}01^{m}57^{s}.75

2002 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green